OpenStudy (anonymous):

without dividing show that 8 is the factor of 256

OpenStudy (amistre64):

use them as 2^k numbers

OpenStudy (amistre64):

show that 8 = 2^k, and 256 = 2^(kn)

OpenStudy (amistre64):

err, 2^(k+n) :)

OpenStudy (anonymous):

8=2^3 and 256=2^8 now???

OpenStudy (amistre64):

8 = 3+?

OpenStudy (amistre64):

\[a^{k+n}=a^k~a^n\]

OpenStudy (amistre64):

since 8 = 2^3; 256 = 2^(3+5) =2^3 2^5 = 8 2^5

OpenStudy (anonymous):

ok get it...thanks

OpenStudy (amistre64):

youre welcome